OpenStudy (anonymous):

A. anumber increased by 13 results in -100. B.find the number.

OpenStudy (kainui):

A number is your variable, so we'll call it x. Increased by 13 means +13 right? "Results in" is the same thing as "equal to" so we can replace that with =. It says it should be equal to -100, right? So lets try to make an equation out of that.

OpenStudy (kainui):

Try to make it into an equation, and I'll help you figure it out.

OpenStudy (kainui):

Exactly! Now all you have to do for part B is solve for x, do you know how to do that?

OpenStudy (kainui):

Alright, so to get x, you need it to be all by itself on one side. Since there's an equals sign, you know that it's an equality. All that means is that everything on the left side is equal to everything on the right side. So in order to keep things equal, you just have to do the same thing to both sides any time you do something to it. So for x+13=-100 you want to get x by itself, right? So what you can do is subtract 13 from the left side so that it's just x. But, since it's an equality like I said up above, you also have to subtract 13 from the other side to keep things equal. Think you can figure it out from there?

OpenStudy (anonymous):

would the answer be 13/100

OpenStudy (anonymous):

can I please get a response?

OpenStudy (kainui):

Didn't notice you responded, got distracted! The answer would look like this: x+13 = -100 Now if you subtract 13 from the left side x+13-13 you have just x. but since it's an equality you have to subtract 13 from -100 aswell. -100-13=-113 So your answer will be x=-113 x+13=-100 -13 -13 x=-113

OpenStudy (kainui):

No problem, glad I could help.